What is the real and nominal value of a 2.5% raise on a $72,000 salary with...

  1. What is the real and nominal value of a 2.5% raise on a $72,000 salary with a .75% increase in inflation?
  2. You are the AD of a high school and have to replace your field in 10 years. You anticipate it will cost $500,000. Will you be able to fund it if you put $25,000 annually into an account that earns 5% interest?
  3. What is the present value of a sponsorship paying 75,000 a year for 10 years at an 8% interest rate?
  4. Is this a better investment than a sponsorship paying 65,000 a year for 12 years at an 8% interest rate?
Answer #1

Hi,

As per policy we will solve only one question here.

Here Base salary = $72,000

nominal raise = 2.5%

hence nominal value = 72000*(1+2.5%) = $73,800

Real rate = (1+nominal rate)/(1+inflation)-1

= (1+2.5%)/(1+0.75%)-1

= 1.025/1.0075-1

= 0.01737 = 1.737%

hence real value = 72000*(1+1.737%) = $73,250.62

Thanks